# Designing Consensus-Based Distributed Filtering over Directed Graphs

### v1

@@ Abstract
This paper proposes a novel consensus-on-only-measurement distributed filter over directed graphs under the collectively observability condition. First, the distributed filter structure is designed with an augmented leader-following measurement fusion strategy. Subsequently, two parameter design methods are presented, and the consensus gain parameter is devised utilizing local information exclusively rather than global information. Additionally, the lower bound of the fusion step is derived to guarantee a uniformly upper bound of the estimation error covariance. Moreover, the lower bounds of the convergence rates of the steady-state performance gap between the proposed algorithm and the centralized filter are provided with the fusion step approaching infinity. The analysis demonstrates that the convergence rate is, at a minimum, as rapid as exponential convergence under the spectral norm condition of the communication graph. The transient performance is also analyzed with the fusion step tending to infinity. The inherent trade-off between the communication cost and the filtering performance is revealed from the analysis of the steady-state performance and the transient performance. Finally, the theoretical results are substantiated through the validation of two simulation examples.
keywords:
Distributed filtering; Consensus; Sensor networks; Performance analysis; Directed graph.
1 Introduction
@@
In the preceding two decades, investigations into sensor networks have occupied a prominent position in the realm of systems and control, owing to its extensive applications encompassing health care monitoring[16, 1], environmental sensing[8, 12], and collaborative mapping [11, 25]. In order to meet the demands for the network reliability and alleviate the computational and communicative burdens on energy-constrained sensors, the implementation of distributed state estimation serves as a good solution to these challenges. The objective of distributed state estimation is to estimate the state of the target system for each sensor in sensor networks by utilizing information from local neighbors, including local measurements, local state estimates, or related data. Founded on various dynamical systems and constraint conditions, a wealth of research achievements regarding the distributed state estimation have emerged.
The evolution of consensus theory within the domain of multi-agent systems has imparted profound insights into the information fusion for distributed state estimation [24, 15, 29, 20]. Several reputable consensus-based approaches to distributed state estimation have been advanced [22, 23, 7, 4, 6, 30]. In [21, 22], consensus filters, including low-pass, high-pass, and band-pass filters, were employed to fuse the covariances, measurements, or state estimates for both continuous-time and discrete-time systems. In [23], consensus terms were designed in the filter structure, and the optimal gain was derived following with a Lyapunov-based stability analysis. Diffusion strategies, where each sensor transmitted the information to its neighbors, were proposed for distributed Kalman filtering and smoothing in [7]. In [4], consensus of the normalized geometric mean of the probability density functions, called the Kullback-Leibler average, was proposed, and this fusion rule aligned with the concepts of the generalized covariance intersection. In [6], the consensus algorithms on measurements and information were combined to preserve the positive attributes of both approaches, thereby enhancing energy efficiency while diminishing the conservative assumption regarding noise correlation. A distributed state estimation filter by using consensus-based information fusion strategies for continuous-time systems with correlated measurement noise was proposed in [10].
The applicability of distributed algorithms is influenced by the constraints imposed on the communication topologies. There usually exist two key assumptions: undirected connected graphs [22, 9, 28, 10, 19] and directed and strongly connected graphs [4, 6, 30]. In general, directed graphs can model more intricate communication topologies compared to undirected graphs. For directed graphs, the works [4, 6] stated that the consensus matrix needed to be primitive and doubly stochastic to achieve an average consensus, with the introduction of Metropolis weights under the assumption that the graph is undirected. However, challenges arise when applying these principles to directed graphs. Hence, this paper focuses on designing algorithms and their corresponding parameters over directed graphs characterized by an adjacent matrix with elements 00 and $1$.
The classical distributed Kalman filter structure is characterized by the iterations of vectors and covariance matrices. To avoid the computation and transmission of numerous matrices, a consensus-on-only-measurement strategy is considered in this paper. Furthermore, the detailed performance of the proposed distributed filter is analyzed. There exist various performance analysis methods for distributed filters. The Lyapunov’s second method for stability was utilized in many works [23, 5, 4, 6] by constructing the Lyapunov function and conducting stability analysis. Recently, the performance of the distributed filter was analyzed by investigating the difference between the solutions of a modified discrete-time algebraic Riccati equation and a discrete-time Lyapunov equation in [26]. For distributed filters, the performance gap between the distributed filter and the centralized filter is a significant concern, and this paper also focuses on this aspect.
Motivated by the aforementioned observations, this paper aims to propose a consensus-on-only-measurement distributed filter over directed graphs, and furnish comprehensive performance analysis of the proposed distributed filter. The main contributions of this paper are summarized below:
@@
    A novel consensus-on-only-measurement distributed filter (COMDF) over directed graphs is proposed for discrete-time systems under the collectively observability condition, and an augmented leader-following measurement fusion strategy is presented to estimate the other sensors’ measurements. Transmitting only measurements eliminates the need for the computation and transmission of numerous matrices. Additionally, the properties of the measurement estimate error and the state estimate error are presented (Proposition 1 and 2).
@@
    The parameter design methods are provided, including a distributed design method for each sensor’s consensus gain and the establishment of a lower bound of the fusion step (Theorem 1). Compared to the unified design method [10], the proposed distributed design method exclusively relies on local information, facilitating the distributed usage without the need for global information. The lower bound of the fusion step illuminates the necessary conditions to guarantee the uniformly upper bound of the estimation error covariance.
@@
    The impact of the fusion step on the steady-state performance of COMDF is analyzed under the conditions involving the spectral radius and the spectral norm of the communication matrix. The lower bounds of the convergence rates of the steady-state performance gap between COMDF and the centralized filter are provided with the fusion step approaching infinity. Particularly, under the spectral norm condition, the convergence rate is at least as fast as exponential convergence (Theorem 3). Additionally, the transient performance is also discussed with the fusion step tending to infinity (Theorem 4).
@@
The remainder of this paper is organized as follows. In Section 2, some preliminaries, including graph concepts, system models, and some useful lemmas, and the problem formulation are provided. In Section 3, a consensus-based distributed filter is proposed over directed graphs, and two parameter design methods are given. In Section 4, the steady-state and the transient performances of the proposed filter are analyzed with the increasing fusion step. In Section 5, two numerical examples are provided to validate the effectiveness of the obtained results. In Section 6, conclusions are drawn.
@@
Notations: Throughout this paper, define $\mathcal{R}^{n}$ as the sets of $n$-dimensional real vectors and $\mathcal{R}^{n\times n}$ as $n\times n$-dimensional real matrices. For a matrix $A\in\mathcal{R}^{n\times n}$, let $A^{-1}$ and $A^{T}$ represent its inverse and transpose, respectively, $\|A\|_{2}$ is the spectral norm, $\rho(A)$ is the spectral radius, and $[A]_{ij}$ denotes the $(i,j)$-th element of the matrix $A$. Notation $\otimes$ represents the Kronecker product. The matrix inequalities $A>B$ and $A\geq B$ signify that $A-B$ is positive definite and positive semi-definite, respectively. $E\{x\}$ denotes the expectation of the random variable $x$.
2 Preliminaries and Problem Statement
2.1 Graph Theory
The communication topology ${\mathcal{G}}(\mathcal{V},\mathcal{E})$ is utilized to illustrate the nodes and the communication links within a sensor network, where the node set $\mathcal{V}=\{1,2,\ldots,N\}$ and the edge set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. For $i,j\in\mathcal{V}$, $(j,i)$ signifies that node $j$ can transmit information to node $i$, and node $j$ is called a neighbor of node $i$. The neighbor set of node $i$ is denoted as $\mathcal{N}_{i}=\{j|(j,i)\in\mathcal{V}\}$, and $|\mathcal{N}_{i}|$ is the cardinality of the neighbors of node $i$. The adjacent matrix is $S=[a_{ij}]_{N\times N}$, where $a_{ij}=1$ if $(j,i)\in\mathcal{E}$ and $a_{ij}=0$ otherwise. Let $D=\text{diag}\{|\mathcal{N}_{1}|,\ldots,|\mathcal{N}_{N}|\}$, and the Laplacian matrix is defined as $\mathcal{L}=D-S$. The edge $(i,j)$ is undirected if $(i,j)\in\mathcal{E}$ implies $(j,i)\in\mathcal{E}$. The communication graph is termed undirected if every edge is undirected. The graph $G$ contains a directed path from node $i_{1}$ to node $i_{m}$, if there exists a sequence of connected edges $(i_{k},i_{k+1}),k=1,\ldots,m-1$. The communication graph is called strongly connected, if there exists a path between any pair of distinct nodes.
Assumption 1.
The communication graph is directed and strongly connected.
Remark 1.
Assumption 1 can be extended to a jointly connected switching topology by applying the corresponding definition of communication networks [27, 17, 18]. The proposed parameter design methods are applicable in these situations as well. For simplicity, this paper focuses on the directed and strongly connected topology assumption.
2.2 System Model
@@
$y_{i,k}=C_{i}x_{k}+\nu_{i,k},~{}~{}~{}i=1,2,...,N,$
where $k$ is the discrete-time index, $i\in\mathcal{V}$ is the $i$-th sensor of the network, $x_{k}\in\mathcal{R}^{n}$ is the system state vector, $y_{i,k}\in\mathcal{R}^{r_{i}}$ is the measurement vector taken by sensor $i$, $A\in\mathcal{R}^{n\times n}$ is the state transition matrix, $C_{i}\in\mathcal{R}^{r_{i}\times n}$ is the observation matrix of sensor $i$, and $\omega_{k}\in\mathcal{R}^{n}$ and $\nu_{i,k}\in\mathcal{R}^{r_{i}}$ are the zero-mean Gaussian noise with the covariances $Q_{k}\in\mathcal{R}^{n\times n}$ and $R_{i,k}\in\mathcal{R}^{r_{i}\times r_{i}}$, respectively. The noise sequences $\{\omega_{k}\}^{\infty}_{k=0}$ and $\{\nu_{i,k}\}^{\infty,N}_{k=0,i=1}$ are mutually uncorrelated. For the whole network, $C=[C^{T}_{1},\ldots,C^{T}_{N}]^{T}\in\mathcal{R}^{r\times n}$ is the augmented observation matrix, and $R=\text{diag}\{R_{1},\ldots,R_{N}\}\in\mathcal{R}^{r\times r}$ is the augmented measurement noise covariance matrix, where $r=\sum^{N}_{i=1}r_{i}$.
Assumption 2.
@@
2.3 Some Useful Lemmas
Definition 1.
[13] For a matrix $M\in\mathcal{R}^{n\times n}$, $M$ is said to be irreducibly diagonally dominant if
@@ 2.3 Some Useful Lemmas
    $M$ is irreducible,
@@
    $M$ is diagonally dominant, i.e., $|m_{ii}|\geq R^{{}^{\prime}}_{i}(M)$ for all $i=1,\ldots,n$,
  3. 3.
    There is an $i\in\{1,\ldots,n\}$ such that $|m_{ii}|>R^{{}^{\prime}}_{i}(M)$,
where $R^{{}^{\prime}}_{i}(M)=\sum_{j\neq i}|m_{ij}|$.
Lemma 1.
[13] Let the matrix $M$ be irreducibly diagonally dominant. Then,
  1. 1.
    $M$ is nonsingular,
  2. 2.
    If $M$ is Hermitian and every main diagonal entry is positive, $M$ is positive definite.
Lemma 2.
Lemma 3.
[13] For a nonnegative matrix $M=[m_{ij}]$, it holds
$\text{min}_{1\leq i\leq n}\sum^{n}_{j=1}m_{ij}\leq\rho(M)\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}m_{ij}.$ (2)
Lemma 4.
[26] For any matrix $M\in\mathcal{R}^{n\times n}$, it holds
$\|M^{k}\|_{2}\leq\sqrt{n}\sum^{n-1}_{j=0}\binom{n-1}{j}\binom{k}{j}\|M\|^{j}_{2}\rho(M)^{k-j},$ (3)
where $\binom{m}{n}$ is the combinatorial number with $\binom{k}{j}=0$ for $j>k$.
2.4 Problem Statement
  1. 1.
    Develop a distributed filter algorithm over directed graphs, utilizing only measurements. Explore parameter design methods to guarantee the convergence and stability of the distributed filter.
  2. 2.
    Evaluate how the fusion step influences the steady-state and transient performance. Find out the performance gap between the proposed distributed filter and the centralized filter.
3 Distributed Filter
In this section, a consensus-on-only-measurement distributed filter (COMDF) is proposed. Subsequently, the state estimate error and the measurement estimate error are defined, and their properties are analyzed. Furthermore, two parameter design methods are provided to guarantee the stability and convergence of the proposed distributed filter.
3.1 Design of the Distributed Filter
This subsection proposes the consensus-on-only-measurement distributed filter, and an augmented leader-following measurement fusion strategy is designed to estimate the neighbors’ measurements.
The state estimator structure of the target system (1) for sensor $i$ is designed as
$\hat{x}_{i,k|k-1}=A\hat{x}_{i,k-1|k-1},{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}$ (4)
$\hat{x}_{i,k|k}=\hat{x}_{i,k|k-1}+K(z^{(l)}_{i,k}-C\hat{x}_{i,k|k-1}),$ (5)
where $z^{(l)}_{i,k}$ represents the sensor $i$’s estimate of measurements from other sensors at the $l$-th consensus step, the gain matrix $K$ is given by
$K=PC^{T}(CPC^{T}+R)^{-1},$ (6)
and $P$ is determined by solving the discrete algebraic Riccati equation
$P=APA^{T}+Q-APC^{T}(CPC^{T}+R)^{-1}CPA^{T}.$ (7)
The sensor $i$’s estimate of the sensor $j$’s measurement at the $l$-th fusion step, denoted as $z^{(l)}_{ij,k}$, is computed as
$z^{(0)}_{ij,k}=C_{j}\hat{x}_{i,k|k-1},{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}$ (8)
$z^{(l)}_{ij,k}$ $=z^{(l-1)}_{ij,k}-\mu_{ij}\Big{[}\sum^{N}_{l=1}a_{il}(z^{(l-1)}_{ij,k}-z^{(l-1)}_{lj,k})$ (9)
@@ 3.1 Design of the Distributed Filter
where the consensus gain $\mu_{ij}$, designed later, is a positive constant, and $a_{ij}$ is the element of the adjacent matrix of the corresponding communication topology. By denoting the augmented vectors as $z^{(l)}_{i,k}=[(z^{(l)}_{i1,k})^{T},\ldots,(z^{(l)}_{iN,k})^{T}]^{T}$ and $y_{k}=[y^{T}_{1,k},\ldots,y^{T}_{N,k}]^{T}$, $z^{(l)}_{i,k}$ can be expressed as
$z^{(0)}_{i,k}=C\hat{x}_{i,k|k-1},{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}{}$ (10)
$~{}~{}~{}~{}z^{(l)}_{i,k}$ $=z^{(l-1)}_{i,k}-\Lambda_{i}\Big{[}\sum^{N}_{l=1}a_{il}(z^{(l-1)}_{i,k}-z^{(l-1)}_{l,k})$ (11)
@@
Now, the distributed estimator is constructed based on (4), (5), and (11).
Remark 2.
An augmented leader-following measurement information fusion strategy is designed in the algorithm, with only the transmission of measurement estimates in the sensor networks. The gain matrix $K$ can be precomputed, where the matrix $C$ and $R$ can be obtained by using a similar consensus algorithm as presented in (8) and (9). Two methods are presented here to obtain the corresponding matrices: one involves vectorizing the matrix and using the consensus strategy (9), while the other employs a matrix consensus strategy by replacing the vectors in (9) with the corresponding matrices.
Remark 3.
The proposed distributed estimator has the following advantages. First, it only needs the addition and substraction of the measurement vectors, thereby avoiding the computation and transmission of plenties of matrices. Second, the proposed distributed filter is asymptotically optimal as the fusion step $l$ tends to infinity. Third, it exhibits a stronger privacy protection performance with only the transmission of the measurements rather than the state estimates like [23, 14]. Fourth, the algorithm can be employed in directed graphs, and the parameters can be designed by only using local neighbors’ information.
Remark 4.
The local gain matrix $K_{i}$, substituting for $K$ in (5), can also be utilized in the algorithm. By using local information regarding the measurement matrices and the measurement noise covariances, $K_{i}$ can be obtained and utilized. Under the local observability condition, this algorithm also works. To better present the properties of the proposed filter, the global $K$ is considered. There exist two parameters $\mu_{ij}$ and $l$, and the parameter design methods are introduced later.
3.2 Two Estimation Errors
Define the state estimation error and the measurement estimation error as
$e_{i,k|k}=\hat{x}_{i,k|k}-x_{k},$
and
$\varepsilon^{(l)}_{i,k}=z^{(l)}_{i,k}-y_{k},~{}~{}~{}~{}$
respectively. Denote the augmented vectors as $e_{k}=[e^{T}_{1,k},\ldots,e^{T}_{N,k}]^{T}$, $\varepsilon^{(l)}_{k}=[(\varepsilon^{(l)}_{1,k})^{T},\ldots,(\varepsilon^{(l)}_{N,k})^{T}]^{T}$, $z^{(l)}_{k}=[(z^{(l)}_{1,k})^{T},\ldots,(z^{(l)}_{N,k})^{T}]^{T}$, and $\nu_{k}=[\nu^{T}_{1,k},\ldots,\nu^{T}_{N,k}]^{T}$. Next, the statistical properties of the measurement estimation error $\varepsilon^{(l)}_{k}$ and the state estimation error $e_{i,k|k}$ are derived.
Proposition 1.
For the measurement estimation error $\varepsilon^{(l)}_{k}$ and its estimation error covariance $P_{\varepsilon,k}=E\{\varepsilon^{(l)}_{k}(\varepsilon^{(l)}_{k})^{T}\}$, the following results hold

### v2

# Distributed State Estimation for Discrete-Time Linear Systems over Directed Graphs: A Measurement Perspective

@@ Abstract
This paper proposes a novel consensus-based distributed filter over directed graphs under the collectively observability condition. The distributed filter is designed using an augmented leader-following information fusion strategy, and the gain parameter is determined exclusively using local information. Additionally, the lower bound of the fusion step number is derived to ensure that the estimation error covariance remains uniformly upper-bounded. Furthermore, the lower bounds for the convergence rates of the steady-state performance gap between the proposed filter and the centralized filter are provided as the fusion step number approaches infinity. The analysis demonstrates that the convergence rate is at least as fast as exponential convergence, provided the communication topology satisfies the spectral norm condition. Finally, the theoretical results are validated through two simulation examples.
I Introduction
@@
In the preceding two decades, investigations into sensor networks have occupied a prominent position in the realm of systems and control, owing to its extensive applications encompassing health care monitoring [13, 1], environmental sensing [7, 10], and collaborative mapping [9, 18]. In order to meet the demands for the network reliability and alleviate the computational and communicative burdens on energy-constrained sensors, the implementation of distributed state estimation serves as a good solution to these challenges. The objective of distributed state estimation is to estimate the state of the target system for each sensor in sensor networks.
Building on various dynamical systems and constraint conditions, a wealth of research has emerged in the field of distributed state estimation. The evolution of consensus theory within the domain of multi-agent systems has imparted profound insights into the information fusion for distributed state estimation [15, 12, 22]. Several reputable consensus-based approaches to distributed state estimation have been proposed [16, 23, 3, 5, 17, 4, 8]. Depending on the filter structures, these filters fuse different types of information from local neighbors, including local state estimates [5, 23, 3], intermediate quantities [17, 4], or measurements [8]. For the fusion of local state estimates, Battistelli et al. [5] introduced a distributed filter based on the covariance intersection method, which combines estimates from different sensors. In another approach, Battilotti et al. [3] developed a distributed filter that fuses the state estimation deviations among neighboring sensors. Regarding the fusion of intermediate quantities, Olfati-Saber [17] designed a Kalman-consensus filter using intermediate quantities to approximate the centralized filter, with a Lyapunov-based stability analysis. Additionally, Battistelli et al. [4] presented a distributed filter that fuses both state estimates and intermediate quantities, thereby preserving the benefits of both approaches. For the fusion of local measurements, Duan et al. [8] proposed a distributed state estimation filter for continuous-time systems with correlated measurement noise.
However, compared to existing literature, the research on the fusion of local measurements has not been thoroughly investigated for discrete-time linear systems due to the following challenges:
@@
    Designing the filter structure and information fusion strategies using local measurements is challenging while ensuring the filter performance.
@@
    Existing gain design methods for directed graphs [5, 4] face complexities in applications, and developing an effective distributed design method is crucial.
@@
    It is necessary to explore the performance gap between the proposed filter and the centralized filter, with a particular focus on the convergence rate regarding the fusion step number.
Motivated by the above observations, this paper aims to propose a novel consensus-based distributed filter over directed graphs using a dynamic averaging algorithm, and provides a comprehensive performance analysis of the proposed distributed filter. The main contributions of this paper are summarized below:
  1. 1.
    We propose a novel consensus-based distributed filter using an augmented leader-following information fusion strategy over directed graphs under the collectively observability condition. Compared to the existing literature [16, 23, 3, 5, 17, 4, 8], this filter provides a novel filter structure that has been rarely studied.
  2. 2.
    Under the coupling between the measurement estimation error and the state estimation error, we provide gain design methods that rely solely on local information. Additionally, we derive the lower bound for the fusion step number (Theorem 1).
  3. 3.
    We analyze the impact of the fusion step number on the steady-state performance. Specifically, we provide the lower bounds on the convergence rates of the steady-state performance gap between the proposed distributed filter and the centralized filter as the fusion step approaches infinity (Theorem 3).
@@
The remainder of this paper is organized as follows. Section II presents the preliminaries and problem formulation. Section III proposes a consensus-based distributed filter over directed graphs, and provides two parameter design methods. Section IV analyzes the steady-state performance of the proposed filter as the fusion step number increases. Section V presents two numerical examples to validate the effectiveness of the results. Finally, Section VI concludes the paper.
@@
Notations: For a matrix $A\in\mathcal{R}^{n\times n}$, $\|A\|_{2}$ is the spectral norm, $\rho(A)$ is the spectral radius, and $A]_{ij}$ denotes the $(i,j)$-th element of the matrix $A$. Symbol $\otimes$ represents the Kronecker product. The matrix inequalities $A>B$ and $A\geq B$ signify that ABA-B is positive definite and positive semi-definite, respectively.
II Preliminaries and Problem Statement
II-A Some Useful Lemmas
Definition 1
[11] For a matrix $\in\mathcal{R}^{n\times n}, MM$ is said to be irreducibly diagonally dominant if 1) $M$ is irreducible, 2) $M$ is diagonally dominant, i.e., $|m_{ii}|\geq R^{s}_{i}(M)$ for all $i=1,\ldots,n$, 3) There is an $i\in\{1,\ldots,n\}$ such that $|m_{ii}|>R^{s}_{i}(M)$, where $R^{s}_{i}(M)=\sum_{j\neq i}|m_{ij}|$.
Lemma 1
[11] Let the matrix $M$ be irreducibly diagonally dominant. Then, 1) $M$ is nonsingular, 2) If $M$ is Hermitian and every main diagonal entry is positive, $M$ is positive definite.
Lemma 2
Lemma 3
[11] For a nonnegative matrix $M=[m_{ij}]$, it holds $\text{min}_{1\leq i\leq n}\sum^{n}_{j=1}m_{ij}\leq\rho(M)\leq\text{max}_{1\leq i\leq n}\sum^{n}_{j=1}m_{ij}.$
Lemma 4
[19] For any matrix $M\in\mathcal{R}^{n\times n}$, it holds $\|M^{k}\|_{2}\leq\sqrt{n}\sum^{n-1}_{j=0}\binom{n-1}{j}\binom{k}{j}\|M\|^{j}_{2}\rho(M)^{k-j},$ where $\binom{m}{n}$ is the combinatorial number with $\binom{k}{j}=0$ for $j>k$.
II-B System Model
@@
$y_{i,k}=C_{i}x_{k}+\nu_{i,k},~~~i=1,2,.,N,$
where $k$ represents the discrete-time index, $i$ represents the $i$-th sensor in the network, $x_{k}\in\mathcal{R}^{n}$ is the system state vector, $y_{i,k}\in\mathcal{R}^{r_{i}}$ is the measurement vector of sensor $i$, $A\in\mathcal{R}^{n\times n}$ is the state transition matrix, $C_{i}\in\mathcal{R}^{r_{i}\times n}$ is the observation matrix, and $\omega_{k}\in\mathcal{R}^{n}$ and $\nu_{i,k}\in\mathcal{R}^{r_{i}}$ are zero-mean Gaussian noise with the covariances $Q_{k}\in\mathcal{R}^{n\times n}$ and $R_{i,k}\in\mathcal{R}^{r_{i}\times r_{i}}$, respectively. The noise sequences $\{\omega_{k}\}^{\infty}_{k=0}$ and $\{\nu_{i,k}\}^{\infty,N}_{k=0,i=1}$ are mutually uncorrelated. Let $C=[C^{T}_{1},\ldots,C^{T}_{N}]^{T}\in\mathcal{R}^{r\times n}$ denote the augmented observation matrix, and $R=\text{diag}\{R_{1},\ldots,R_{N}\}\in\mathcal{R}^{r\times r}$ represent the augmented measurement noise covariance matrix, where $r=\sum^{N}_{i=1}r_{i}$. The communication topology of the sensor network is described by a graph ${\mathcal{G}}(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{1,2,\ldots,N\}$ denotes the sensor set, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the edge set defining the communication links among the sensors. For $i,j\in\mathcal{V}$, the edge $(j,i)$ indicates that sensor $j$ can transmit information to sensor $i$, and sensor $j$ is referred as a neighbor of sensor $i$. The neighbor set of sensor $i$ is denoted as $\mathcal{N}_{i}=\{j|(j,i)\in\mathcal{V}\}$, and $|\mathcal{N}_{i}|$ represents the number of neighbors of sensor $i$. The adjacency matrix is defined as $S=[a_{ij}]_{N\times N}$, where $a_{ij}=1$ if $(j,i)\in\mathcal{E}$ and $a_{ij}=0$ otherwise. Let $D=\text{diag}\{|\mathcal{N}_{1}|,\ldots,|\mathcal{N}_{N}|\}$, and the Laplacian matrix is defined as $\mathcal{L}=D-S$. An edge $(i,j)$ is undirected if $(i,j)\in\mathcal{E}$ implies $(j,i)\in\mathcal{E}$. The communication graph is termed undirected if all edges are undirected. The graph $G$ contains a directed path from sensor $i_{1}$ to sensor $i_{m}$, if there exists a sequence of connected edges $(i_{k},i_{k+1}),k=1,\ldots,m-1$. The communication graph is called strongly connected, if there exists a path between any pair of distinct nodes.
Assumption 1
@@
Assumption 2
The communication graph is directed and strongly connected.
Remark 1
Assumption 1 shows that observability can be achieved by the entire network, even if it is not satisfied by any single sensor. Assumption 2 can be extended to a jointly connected switching topology by applying the corresponding definition of communication networks [21, 14].
II-C Problem Statement
@@
    Develop a distributed filter over directed graphs with a leader-following information fusion strategy using local measurements. Explore parameter design methods to guarantee the filter’s performance.
@@
    Find out the performance gap between the proposed distributed filter and the centralized filter. Evaluate how the fusion step number influences the performance.
III Distributed Filter
In this section, we propose a novel consensus-based distributed filter. Subsequently, we define the state estimation error and the measurement estimation error, and analyze their properties. Furthermore, we provide two parameter design methods to guarantee the filter’s performance.
III-A Design of the Distributed Filter
This subsection proposes a novel distributed state estimator over directed graphs with the consensus-based information fusion strategy.
First, design sensor $i$’s estimate of sensor $j$’s measurement at the $l$-th fusion step, denoted as $z^{(l)}_{ij,k}$, as follows:
$z^{(0)}_{ij,k}=C_{j}A\hat{x}_{i,k-1|k-1},$ (2)
$z^{(l)}_{ij,k}$ $=z^{(l-1)}_{ij,k}-\mu_{ij}\Big[\sum_{t\in\mathcal{N}_{i}}a_{it}(z^{(l-1)}_{ij,k}-z^{(l-1)}_{tj,k})$ (3)
@@ III-A Design of the Distributed Filter
where the consensus gain $\mu_{ij}$, to be designed later, is a positive constant, and $a_{ij}$ is the element of the adjacency matrix corresponding to the communication topology. By denoting the augmented vectors as $z^{(l)}_{i,k}=[(z^{(l)}_{i1,k})^{T},\ldots,(z^{(l)}_{iN,k})^{T}]^{T}$ and $y_{k}=[y^{T}_{1,k},\ldots,y^{T}_{N,k}]^{T}$, $z^{(l)}_{i,k}$ can be expressed as
$z^{(0)}_{i,k}=CA\hat{x}_{i,k-1|k-1},$ (4)
$~~~~z^{(l)}_{i,k}$ $=z^{(l-1)}_{i,k}-\Lambda_{i}\Big[\sum^{N}_{t=1}a_{it}(z^{(l-1)}_{i,k}-z^{(l-1)}_{t,k})~~~~$ (5)
@@
Next, the state estimator structure of the system (1) for sensor $i$ is designed as
$\hat{x}_{i,k|k}=(A-KCA)\hat{x}_{i,k-1|k-1}+Kz^{(l)}_{i,k}$, (6)
where the gain matrix $K$ is given by $K=\bar{P}C^{T}R^{-1},$ $\bar{P}=P-PC^{T}(CPC^{T}+R)^{-1}CP$, and $P$ is determined by solving the discrete algebraic Riccati equation $P=APA^{T}+Q-APC^{T}(CPC^{T}+R)^{-1}CPA^{T}$.
Now, the distributed estimator is constructed based on (6), and its structure is illustrated in Fig. 1.
Figure 1: Illustration of the proposed distributed estimator.
Remark 2
In the algorithm, both the measurement estimates $z^{(l-1)}_{it,k}$ and the raw measurements $y_{i,k}$ are transmitted. Between two consecutive time steps, the information is exchanged $l$ times. Since each node may not obtain sufficient measurements from its neighbors, the measurement estimates, obtained through prediction, compensate for this limitation.
Remark 3
The proposed distributed estimator offers several advantages. First, it complements existing distributed filters through a local measurement prediction fusion mechanism, thereby establishing a general framework for further investigation. Second, the proposed distributed filter demonstrates strong performance by achieving asymptotically optimality as the fusion step $l$ tends to infinity. Third, it can be applied to directed graphs with a simpler parameter design method compared to [5, 4]. Fourth, when direct measurement information is unavailable, the predicted measurements compensate for missing data, thus maintaining estimation performance. Finally, its flexibility in communication steps makes it suitable for extensions to event-triggered or adaptive communication strategies, while the consensus framework can also be extended to support privacy-preserving and secure filtering schemes.
III-B Two Estimation Errors
This subsection computes two estimation errors for further analysis. Define the measurement estimation error and the state estimation error as $\varepsilon^{(l)}_{i,k}=z^{(l)}_{i,k}-y_{k},$ and $e_{i,k|k}=\hat{x}_{i,k|k}-x_{k},$ respectively. Denote the augmented vectors as
$e_{k}=[e^{T}_{1,k},\ldots,e^{T}_{N,k}]^{T},~~~~~~~\varepsilon^{(l)}_{k}=[(\varepsilon^{(l)}_{1,k})^{T},\ldots,(\varepsilon^{(l)}_{N,k})^{T}]^{T},$
$z^{(l)}_{k}=[(z^{(l)}_{1,k})^{T},\ldots,(z^{(l)}_{N,k})^{T}]^{T},~\nu_{k}=[\nu^{T}_{1,k},\ldots,\nu^{T}_{N,k}]^{T}.$
Next, we derive the statistical properties of the measurement estimation error $\varepsilon^{(l)}_{k}$ and the state estimation error $e_{i,k|k}$. Additionally, we define their estimation error covariances as $P_{\varepsilon,k}=E\{\varepsilon^{(l)}_{k}(\varepsilon^{(l)}_{k})^{T}\}$ and $P_{k|k}=E\{e_{k}e^{T}_{k}\}$.
Proposition 1
The following statements hold